Solve
6y^2dx-x(2x^3+y)dy=0
1
Expert's answer
2020-10-25T18:40:34-0400

Given DE is

"6y^2dx-x(2x^3+y)dy=0"

After rearranging the above equation ,we can write


"\frac{dx}{dy}-\frac{x}{6y}=\frac{x^4}{3y^2}\\
\implies -\frac{3\frac{dx}{dy}}{x^4}+\frac{1}{2yx^3}=-\frac{1}{y^2}"

Put

"v=\frac{1}{x^3}\implies \frac{dv}{dy}= -\frac{3\frac{dx}{dy}}{x^4}"

Thus, we get


"\frac{dv}{dy}+\frac{1}{2y}v=-\frac{1}{y^2}"

Hence, Integrating Factor is

"I.F=e^{\int 1/2ydy}=\sqrt{y}"

So it follows


"\frac{d}{dy}(\sqrt{y}v)=-y^{-3/2}\\
\implies \sqrt{y}v=-\int y^{-3/2}dy=2y^{-1/2}+C_1\\
\implies v=C_1y^{-1/2}+2y^{-1}\\
\implies \frac{1}{x^3}=C_1y^{-1/2}+2y^{-1}\\"

After simplification we get,


"y{\left(x \right)} = \frac{3 C_{1} \sqrt{x^{9} \left(9 C_{1}^{2} x^{3} + 8\right)}}{2} + \frac{x^{3} \left(9 C_{1}^{2} x^{3} + 4\right)}{2}\\
y{\left(x \right)} = - \frac{3 C_{1} \sqrt{x^{9} \left(9 C_{1}^{2} x^{3} + 8\right)}}{2} + \frac{x^{3} \left(9 C_{1}^{2} x^{3} + 4\right)}{2}"